OpenStudy (anonymous):

how to do this sum?please say me with steps.. 8^-4*5^-6*6^-2/9^-4

OpenStudy (anonymous):

follow the basic operator priorty rules.... top priority :\[\left( \right)\div \times +-\] solve whichever operator comes at the left most position of the priorty chart. so solve first parenthesis then division then multiplication the addition and then subtractioin

OpenStudy (anonymous):

\[\frac{8^{-4}*5^{-6}*6^{-2}} {(9^{-4})}\] \[=\frac{9^{4}}{8^{4} \times 5^{6} \times 6^{2}}\]